OpenStudy (anonymous):

Okay, here is the officially dumbest question I will ever ask... Can someone help me remember GCF and LCM basics... Oh boy, I feel dumb... :/

OpenStudy (anonymous):

How many times the largest number appears in a number = GCF , as LCM is the opposite and you must find the least times that number appears.

OpenStudy (anonymous):

With exception of division... sorry for my translation. Maybe a more experienced person will write an answer more understandable.

OpenStudy (anonymous):

Pick a couple of numbers and I'll give an example of how I find the LCM.

OpenStudy (dumbcow):

GCF: take each number and break it down into factors, find largest number they have in common \[15 = 5*3\] \[25 = 5*5\] largest number in common is 5, so GCF of 15 and 25 is 5 LCM: take multiples of each number, find the smallest multiple they all have in common \[6 \rightarrow 6,12,18,24...\] \[8 \rightarrow 8,16,24,32...\] smallest multiple is 24, LCM of 6 and 8 is 24

OpenStudy (anonymous):

The LCM and the GCF for two numbers are related. \[\large LCM(a,b) = \frac{|a\cdot b|}{GCF(a,b)}\]

OpenStudy (anonymous):

Okay... ThanKS!!!

OpenStudy (anonymous):

There's another interesting way to find the GCF of two numbers: http://en.wikipedia.org/wiki/Euclidean_algorithm